Calculate the PPP-adjusted GDP for each of the four countries, using the information found in the table below.

Instructions: Round your answers to the nearest dollar.

Country GDP
($)
Price level
(%)
PPP-adjusted GDP
Ona 12,000 5
Rye 12,800 -26
Zolfo 14,000 -9
Avon 23,300 22
Answer #1

PPP-adjusted GDP = GDP*(1/(100%+Price level)

Ona PPP-adjusted GDP = 12000*(1/(100%+5%)) = 12000/(1.07) = $ 11428.57

Rye PPP-adjusted GDP = 12800*(1/(100%-26%)) = 12800/(0.74) = $ 17297.3

Zolfo PPP-adjusted GDP = 14000*(1/(100%-9%)) = 14000/(0.91) = $ 15384.62

Avon PPP-adjusted GDP = 23300*(1/(100%+22%)) = 23300/(1.22) = $ 19098.36